OpenStudy (anonymous):

(8x^6y^-3)^1/3 how do you do this?

OpenStudy (mertsj):

It is the same as this: \[\sqrt[3]\frac{8x^6}{y^3}\]

OpenStudy (anonymous):

how do you solve this tho? it's going to be like 2x^3/?

OpenStudy (anonymous):

multply each exponent by \(\frac{1}{3}\)

OpenStudy (anonymous):

so is this going to be like \[2x^2/y^-1?\]

OpenStudy (mertsj):

\[\frac{2x^2}{y}\]